Time value comparisons of single amounts table: 1 . Your opportunity cost is 11%. Personal Finance Problem In exchange for a $20,000 payment today, a well-known company will allow you to choose one of the alternatives shown in the following a. Find the value today of each alternative. b. Are all the alternatives acceptable-that is, worth at least $20,000 today? c. Which alternative, if any, will you take? a. The present value of Alternative A is S The present value of Alternative B is $ The present value of Alternative C is $ b. Are all the alternatives (Round to the nearest cent.) Round to the nearest cent.) Round to the nearest cent.) is, worth at least $20,000 today? (Select the best answer below.) O Yes O No c. Which alternative, if any, will you take? (Select the best answer below.) O A. Alternative C OB. None OC. Alternative B O D. Alternative A : Data Table (Click on the icon located on the top-right corner of the data table below in order to copy its contents into a spreadsheet.) Alternative Single amount $28,500 at end of 3 years $63,500 at end of 9 years $191,500 at end of 20 years
Answer #1

a)

Present value of alternative A = Future value / ( 1 + r)n

Present value of alternative A = 28,500 / ( 1 + 0.11)3

Present value of alternative A = 28,500 / 1.367631

Present value of alternative A = $20,839.0

Present value of alternative B = Future value / ( 1 + r)n

Present value of alternative B = 63,500 / ( 1 + 0.11)9

Present value of alternative B = 63,500 / 2.558037

Present value of alternative B = $24,823.7

Present value of alternative C = Future value / ( 1 + r)n

Present value of alternative C = 191,500 / ( 1 + 0.11)20

Present value of alternative C = 191,500 / 8.062312

Present value of alternative C = $23,752.5

b)

Yes as all the alternatives are above $20,000

c)

Alternative B as it has the highest present value